OpenStudy (anonymous):

*** Will give you a medal *** Suppose a and b vary inversely, and b = 8 when a = 6. Write a function that models the variation and find b when a = 30.

hartnn (hartnn):

when a and b vary inversely, we can write \(ab=k, \: or \: b=\dfrac{k}{a}\) where k is a constant. so, just plug in b=8 and a=6 in the above and find 'k'

OpenStudy (anonymous):

k=48

hartnn (hartnn):

yes, k= 48 is correct :)

OpenStudy (anonymous):

so then when a=30 b=48/30

hartnn (hartnn):

yes, that is also correct, when a=30 b= 48/30 =...?

OpenStudy (anonymous):

1.6

hartnn (hartnn):

yes, or you can keep it as 8/5